OpenStudy (anonymous):

Solve by completing the square: 4x^2+24x=11 Show steps please. Thank you! :)

OpenStudy (anonymous):

I know, i really cant get it, i'm gettin desperate lol :O

OpenStudy (amistre64):

4x^2 +24x = 11 4(x^2 +6x) = 11 4(x^2 +6x+9) = 11 +4(9) x^2 +6x+9 = (11 +4(9))/4 (x+3)^2 = (11 +4(9))/4 x+3 = +- sqrt((11 +4(9)))/2 x = -3 +- sqrt((11 +4(9)))/2

OpenStudy (amistre64):

x = -3 +- sqrt((11 +36))/2 x = -3 +- sqrt(47)/2

OpenStudy (amistre64):

do you have any questions? just let me know

OpenStudy (anonymous):

\[So the final answer looks like this, correct?x=-3\pm \sqrt{47/2}\]

OpenStudy (anonymous):

* So the final answer looks like this? ^^